# RefineStat: Efficient Exploration for Probabilistic Program Synthesis

### v1

@@ 5.4 Ablation Study
Table 4 shows that parameter validity emerges as the most critical component, its removal results in a substantial drop of 14.5 percentage points in compilation success. Nonetheless, all checks contribute meaningfully to overall performance: omitting distribution validity or parse-ability leads to consistent declines in compilation rates (−9% and −5.5%, respectively), underscoring the complementary role these validations play in ensuring soundness of generated programs.
@@
Token Efficiency Analysis. To evaluate the computational cost associated with our framework, we measure the number of tokens consumed in generating the best-performing program under Itergen, Baseline (Unconstrained generation), RefineStat without Refinement Loop (RefineStat w/o RL), and RefineStat using Meta-LLama-3-8B. The cumulative token usage across all ten runs is recorded, and this process is repeated across five independent batches (each using a new set of seeds), from which we report the mean and standard deviation.
In contrast, RefineStat is run five times using five independent seeds, and we track the total number of tokens consumed throughout the entire refinement process in each run.
Dataset Itergen Baseline RefineStat w/o RL RefineStat Token
Mean Std Mean Std Mean Std Mean Std Ratio
Eight Schools $98.8$ $5.3$ $606.8$ $386.3$ $147.6$ $36.7$ $1503.0$ $409.5$ $2.5$
Dugongs $209.3$ $225.9$ $747.4$ $495.0$ $294.4$ $86.3$ $419.0$ $87.1$ $0.6$
GP $131.3$ $8.9$ $663.6$ $410.9$ $155.0$ $21.2$ $1660.0$ $253.5$ $2.5$
Peregrine $132.5$ $21.4$ $884.0$ $538.9$ $180.8$ $103.6$ $1740.0$ $418.0$ $2.0$
Surgical $97.3$ $4.3$ $624.4$ $260.6$ $113.2$ $9.2$ $1275.0$ $1348.1$ $2.0$
Average Token Ratio 1.9x
5.5 Memorization Effect
While canonical textbook datasets such as Eight Schools and Radon are valuable for reproducible benchmarking and comparison to prior work (e.g., Li et al. (2024); PyMC3 implementations), there remains a concern that a model could rely on memorization rather than genuine synthesis when such datasets are well known. Prior work, including BoxLM, has also included dataset descriptions directly in prompts (see Figure 1 in Li et al., 2024).
To assess the extent to which RefineStat depends on memorization, we performed two controlled prompt modifications on all datasets using Meta-Llama 3-8B:
@@ 5.5 Memorization Effect
6 Related Works
@@
Researchers have developed several probabilistic techniques for discovering model structures from observed data, including Bayesian networks (Mansinghka et al., 2006; Lowd and Domingos, 2012), matrix-composition models (Grosse et al., 2012), Markov networks (Gogate et al., 2010), and deep probabilistic models (Gens and Pedro, 2013). Probabilistic programming languages (PPLs) (Salvatier et al., 2016; Phan et al., 2019; Carpenter et al., 2017a) provide ways to flexibly represent these models as programs and support generic inference methods. While these approaches have shown promise, they often require both significant domain expertise and specialized API and/or programming language expertise to implement effectively. Given the recent advances in large language models’ (LLMs) code generation capabilities, synthesizing probabilistic programs using LLMs offers a potentially more accessible and flexible approach to automated model discovery. Therefore, our work RefineStat addresses this opportunity by advancing the SOTA in LLM-based probabilistic program synthesis.
@@
Probabilistic Program Synthesis: There have been many approaches for deterministic program synthesis, which is recently been dominated by LLM-based approaches. Several works (Nori et al., 2015; Saad et al., 2019; Gerasimou et al., 2015; Češka et al., 2019; Ellis et al., 2015) have attempted to synthesize probabilistic programs using classical machine learning or symbolic methods. PSKETCH Nori et al. (2015) applies the notion of sketching for probabilistic program synthesis. Gerasimou et al. (2015) uses a multiobjective optimization genetic algorithm for probabilistic model generation. Saad et al. (2019) presents a technique for automatically constructing probabilistic programs for data analysis, interpretation, and prediction.
@@
Most recently, Li et al. (2024) uses LLM for probabilistic program synthesis. The key idea is to use two interacting LLMs: a generator LLM and a critic LLM that improve the synthesized program iteratively. The paper shows that with instances of GPT4 as the generator and critic (closed LLM with likely >>T weights) can find reasonable probabilistic programs from data. As our evaluation shows, RefineStat significantly improves the ability to find programs that fits the data (recall Table 3) and in contrast to Li et al. (2024), runs only a single small open LLM (<8<8B weights), demonstrating the positive impact of constrained decoding.
@@
Program Synthesis with Constrained LLM Decoding: Recent advances in program synthesis have enabled constrained decoding approaches where LLMs generate code while adhering to formal language specifications. These constraints can be partially precomputed and enforced more efficiently for regular (Deutsch et al., 2019; Willard and Louf, 2023; Kuchnik et al., 2023) or context-free (Koo et al., 2024; Ugare et al., 2024b; Dong et al., 2024; Banerjee et al., 2025; Suresh et al., 2025; Firestone et al., 2025) languages, ensuring syntactic correctness. For more dynamic program generation, Poesia et al. (2022) and Ugare et al. (2025) implement error-driven backtracking when the LLM synthesizes invalid code.
[…]
@@ 7 Conclusion and Limitations
Conclusion.
The main contribution of our work is to separate the task of generating probabilistic modeling through PPLs as fragments of priors and likelihood and to construct an LLM-based search procedure that automatically discovers the probabilistic program that satisfies the standard reliability metrics in the Bayesian workflow. We believe that our framework can be extended to enforce arbitrary reliability criteria defined by domain experts for reliable generation in other domains that involve domain-specific languages and plan to explore those in future work.
Limitations.
While our framework incorporates key components of the Bayesian workflow, such as convergence diagnostics and predictive performance metrics, it does not currently include prior-predictive or posterior-predictive checks, which often require manual inspection and domain-specific judgment Gelman et al. (1995). As a result, the reliability assessment is based on a subset of available diagnostics, and in some cases, the reported ELPD values may not fully reflect model adequacy. Moreover, our refinement strategy, while effective in practice, does not guarantee convergence to a fully satisfactory or globally optimal program.
Appendix
@@ Conclusion.
A. Probabilistic Program Metrics
  • $\circ$
    $\widehat{R}$ Statistic
  • $\circ$
    Effective Sample Size (ESS)
  • $\circ$
    Divergent Transitions
  • $\circ$
    Bayesian Fraction of Missing Information (BFMI)
  • $\circ$
    Pareto Shape Parameter
  • $\circ$
    Expected Log Pointwise Predictive Density (ELPD)
B. Error Analysis
  • $\circ$
@@ B. Error Analysis
C. Prompt Design
  • $\circ$
    Prompt Template
  • $\circ$
    Dataset-Specific Templates:
Appendix A Probabilistic Program Metrics
@@
Appendix B Error Analysis
@@
Appendix C Prompt Design
C.1 Eight Schools
Description: A hierarchical model for the 8-schools data.
C.2 Dugongs
Description: A growth model for dugongs with missing data.
C.3 Surgical
C.4 GP
Description: Simulated data from a Poisson GP model.
C.5 Peregrine

### v2

@@ 5.4 Ablation Study
Table 4 shows that parameter validity emerges as the most critical component, its removal results in a substantial drop of 14.5 percentage points in compilation success. All checks contribute meaningfully to overall performance: omitting distribution validity or parse-ability consistently reduce compilation rates (−9% and −5.5%, respectively), underscoring the complementary role these validations play in ensuring soundness of generated programs.
Memorization Effect. A number of studies Dong et al. (2024a); Golchin and Surdeanu (2025; 2024); Li (2023) have highlighted concerns about the memorization effect in large language models, where models may reproduce previously seen content rather than demonstrating genuine synthesis. Kong et al. (2025) addresses this issue by proposing code mutation to reveal potential memorization in program repair tasks. Inspired by this perspective, we introduce dataset and prompt modifications designed to preserve the semantics of the data while changing its presentation. We apply two systematic prompt modifications across all benchmarks when evaluating Meta-Llama 3-8B, with details provided in Appendix H.2. The modified versions match the original RefineStat on reliability, convergence, and predictive metrics. Additionally, PosteriorDB provides limited ground-truth PyMC programs: only the Eight Schools model is available, and it targets an outdated version of PyMC, while the remaining programs are provided in Stan. While these studies are limited in scope, and measuring the memorization effect is still an open problem, they give an indication that RefineStat’s effectiveness may not be the consequence of just memorization.
5.5 Generalizability Across Probabilistic Programming Backends
Although our primary evaluation uses PyMC, the design of RefineStat is not specific to any single probabilistic programming library. To show generalizability of the framework, we apply RefineStat to NumPyro using the same prompting setup and the Qwen-2.5-Coder–7B model.
Across temperatures, RefineStat substantially improves the fraction of programs that successfully compile and execute. Table 5 shows that the run rate more than doubles relative to Standard unconstrained decoding, consistent with the improvements in PyMC3 experiment (Table 1).
Table 5: Run-rate comparison when using NumPyro as the inference backend.
Temp Standard REFINESTAT
0.2 0.17 0.34
0.3 0.15 0.35
0.4 0.15 0.33
Table 6 reports the metrics for NumPyro generated programs across all benchmarks, as those in Table 2 (for PyMC3). For every dataset, RefineStat attains equal or higher reliability than the Standard baseline. When both approaches achieve similar reliability, RefineStat yields comparable or improved predictive performance. Notably, in the GP task, Standard decoding fails to produce any valid model, whereas RefineStat consistently generates executable programs with stable diagnostics. Overall, these results show that RefineStat’s semantic filtering and diagnostic-aware refinement generalize across PPL backends, improving program validity and sampling quality beyond PyMC.
Table 6: Diagnostics and ELPD-LOO results for NumPyro using Qwen-2.5-Coder–7B, demonstrating the generalizability of RefineStat.
[…]
@@ 5.5 Generalizability Across Probabilistic Programming Backends
8 Schools Standard 7.00 0.00 1.00 0.00 1769.00 0.00 0.00 0.00 0.13 0.00 -31.83 0.00
RefineStat 7.00 0.00 1.00 0.00 1850.00 50.00 0.00 0.00 0.11 0.02 -31.10 0.05
Dugongs Standard 7.00 0.00 1.00 0.00 1739.00 0.00 0.00 0.00 0.04 0.00 8.18 0.00
RefineStat 7.00 0.00 1.00 0.00 1747.00 11.55 0.00 0.00 0.04 0.00 8.19 0.01
Peregrine Standard 6.00 0.00 1.00 0.00 3232.50 81.32 0.00 0.00 0.00 0.00 -368.21 0.04
RefineStat 6.67 0.47 1.00 0.00 1658.00 135.60 0.00 0.00 0.11 0.10 -76.25 122.44
Surgical Standard 6.00 0.00 1.00 0.00 1472.00 0.00 0.00 0.00 0.25 0.00 -245.14 0.00
RefineStat 7.00 0.00 1.00 0.00 1628.00 0.00 0.00 0.00 0.08 0.00 -50.37 0.00
GP Standard
RefineStat 6.67 0.47 1.00 0.00 2344.33 115.46 0.00 0.00 0.09 0.13 -83.07 82.70
6 Related Work
@@
Researchers have presented probabilistic techniques for discovering model structures from observed data, including Bayesian networks (Mansinghka et al., 2006; Lowd and Domingos, 2012), matrix-composition models (Grosse et al., 2012), Markov networks (Gogate et al., 2010), and deep probabilistic models (Gens and Domingos, 2013). Probabilistic programming languages (PPLs) flexibly represent these models as programs but demand substantial domain and API expertise. Recent advances in LLM code generation make synthesizing probabilistic programs more accessible. RefineStat leverages this opportunity to advance the state of the art in LLM-based probabilistic program synthesis.
@@
Probabilistic Program Synthesis: There have been many approaches for deterministic program synthesis, which is recently been dominated by LLM-based approaches. Several works (Nori et al., 2015; Saad et al., 2019; Gerasimou et al., 2015; Češka et al., 2019; Ellis et al., 2015) synthesize probabilistic programs using classical machine learning or symbolic methods. Gerasimou et al. (2015) uses a genetic algorithm for probabilistic model generation. Saad et al. (2019) presents techniques to automatically construct probabilistic programs using Bayesian inference over DSLs defined via probabilistic grammars, enabling qualitative structure discovery and quantitative prediction.
@@
Most recently, Li et al. (2024) uses LLM for probabilistic program synthesis. The paper shows that with instances of GPT4 as the generator and critic (closed LLM) can find reasonable probabilistic programs from data. As our evaluation shows, RefineStat significantly improves the ability to find programs that fits the data (recall Table 3) and in contrast to Li et al. (2024), runs only a single small open LLM (<8<8B weights), demonstrating the benefits of constrained decoding.
@@
Program Synthesis with Constrained LLM Decoding: Recent advances in program synthesis have enabled constrained decoding approaches where LLMs generate code while adhering to formal language specifications. These constraints can be partially precomputed and enforced more efficiently for regular (Deutsch et al., 2019; Willard and Louf, 2023; Kuchnik et al., 2023) or context-free (Koo et al., 2024; Ugare et al., 2024b; Dong et al., 2024b; Banerjee et al., 2025; Suresh et al., 2025; Firestone et al., 2025) languages, ensuring syntactic correctness. Recent grammar-constrained and type-constrained approaches rely on a prefix property (Mündler et al., 2025), where each partial program can be incrementally validated and completed into a syntactic and/or well-typed program. In contrast, RefineStat targets properties that cannot be prefix-checked, since statistical reliability requires full posterior inference rather than purely static validation. For more dynamic program generation, Poesia et al. (2022) and Ugare et al. (2025) implement error-driven backtracking.
Recently, several works have explored probabilistic inference/programming in LM-constrained generation. Loula et al. (2025) guide generation with potential scores and grammar rules, constraining token emission but not model correctness. In contrast, RefineStat generates probabilistic programs, enforces PPL checks during decoding, and retains only those passing Bayesian diagnostics. In Grand et al. (2025), a Planner writes an inference plan that LMs execute to satisfy constraints. RefineStat instead directly writes the probabilistic model and focuses on the quality of the posterior. Ahmed et al. (2025) adjusts next-token probabilities using a verifier so text matches high-level attributes. RefineStat aims for statistical validity, enforcing PPL semantics, and selecting the final program.
@@ 7 Conclusion and Limitations
Conclusion. The main contribution of our work is to separate the task of generating probabilistic modeling through PPLs as fragments of priors and likelihood and to construct an LLM-based search procedure that automatically discovers the probabilistic program that satisfies the standard reliability metrics in the Bayesian workflow. We believe that our framework can be extended to enforce arbitrary reliability criteria defined by domain experts for reliable generation in other domains that involve domain-specific languages and plan to explore those in future work.
Limitations. While our framework incorporates key components of the Bayesian workflow (i.e., convergence diagnostics and predictive performance metrics), it does not include prior-predictive or posterior-predictive checks, which often require manual inspection and domain-specific judgment Gelman et al. (1995). Thus, the reliability judgment is based on a subset of available diagnostics, and the reported ELPD only partially reflect model adequacy in some cases. Further, our refinement strategy is effective in practice but does not guarantee convergence to globally optimal program.
8 Acknowledgments
This research was supported in part by NSF Grants No. CCF-1846354 and CCF-2313028.
9 Reproducibility Statement
We provide the source code of RefineStat as part of the supplementary material that can be used to reproduce our results, and we will also release it as open source. We also provide additional experimental details in the appendix.
Appendix A Appendix
@@
A. Extended Background
  1. 1.
    Language Models
  2. 2.
    Bayesian Workflow
  3. 3.
    Probabilistic Programming Language
B. Probabilistic Program Metrics
  1. 1.
    Basic Terminologies
  2. 2.
    Formal Definitions
C. Example
  1. 1.
    Illustrative Example
  2. 2.
    Examples of Structural Changes Introduced by RefineStat
D. Prompt Design
  1. 1.
    Prompt Template
  2. 2.
    Dataset-Specific Prompt
F. Error Analysis
  1. 1.
@@ F. Error Analysis
G. Ablation Studies
  1. 1.
    Token Efficiency Analysis
  2. 2.
    Memorization Effect
Appendix B Extended Background
B.1 Language Models
B.2 Bayesian Workflow
At its core, a generative probabilistic program often follows a $\mathcal{D}\|\mathcal{P}\|\mathcal{L}$ structure: first fixing the observed data ($\mathcal{D}$), then sampling latent variables under the prior ($\mathcal{P}:p(z)$), and finally conditioning on the data via the likelihood ($\mathcal{L}:p(x\mid z)$).
Modern probabilistic programming languages such as Stan (Carpenter et al., 2017a) and PyMC (Salvatier et al., 2016) streamline this cycle by automating MCMC sampling and providing integrated diagnostics. These include prior predictive checks, convergence measures (e.g., split-$\widehat{R}$, effective sample size, BFMI, divergent NUTS transitions) (Vehtari et al., 2021; Hoffman et al., 2014; Betancourt, 2017), and predictive accuracy metrics such as PSIS-LOO (Vehtari et al., 2017).
These diagnostics guard against model mis‐specification and poor sampling behavior, ensuring that only well-calibrated models are considered for inference or downstream decision-making.
B.3 Probabilistic Programming Language
Probabilistic Programming Languages can be categorized as internal Domain-Specific Languages (DSLs), which embed within a host language and reuse its syntax and tooling (e.g., PyMC Salvatier et al. (2016), NumPyro Phan et al. (2019), Pyro Bingham et al. (2019)), or as external DSLs that define their own syntax and compiler (e.g., Stan Carpenter et al. (2017a)). This representation enables automated model specification while leveraging existing inference algorithms Gordon et al. (2014).
Appendix C Probabilistic Program Metrics
C.1 Basic Terminologies
Before presenting examples for individual diagnostics, we briefly define a few recurring terms that are used throughout:
Chain: An independent run of the sampler that generates a sequence of draws from the posterior distribution. Multiple chains are typically run to verify that results do not depend on initialization.
Convergence: The state in which all chains are sampling from the same region of the posterior distribution. Lack of convergence suggests that the sampler has not fully explored the posterior.
Divergence: A warning issued by the Hamiltonian Monte Carlo (HMC) algorithm indicating that numerical integration failed to follow the posterior geometry accurately. Divergences often signal problematic parameterizations or highly curved posterior regions.
C.2 Formal Definitions
@@
Appendix D Examples
D.1 Illustrative Example
We illustrate RefineStat on a standard Bayesian linear regression. Given a partial program $P$, we want to find a completion $M$ that maximizes the Bayesian reliability score $B(M)$. This example shows how each semantic and diagnostic check prunes or refines candidates.
  1. 1.
    Partial Program $P$
    At timestep $t$, RefineStat generates the following partial program:
    with pm.Model() as linear_model:
    alpha = pm.Normal("alpha", 0, 10)
    beta = pm.Normal("beta", 0, 10)
    sigma = pm.HalfNormal("sigma", 5)
    The LLM must complete the likelihood ($y_{\text{obs}}$).
  2. 2.
    LLM Proposals & Semantic Checks
    Candidate likelihoods are checked against PPL semantics as shown in Table 7:
    Table 7: Semantic filtering of LLM-proposed likelihoods.
    Proposal Check Outcome
    y_obs = pm.ExtNormal("y_obs",
    mu=alpha + beta * x,
    sigma=sigma, observed=y) Distribution validity Reject (hallucinated ExtNormal)
    y_obs = pm.Normal("y_obs",
    mu=alpha + beta * x,
    sd=sigma, observed=y) Parameter validity Reject (deprecated sd vs. sigma)
    y_obs = pm.Normal("y_obs",
    mu=alpha + beta * x,
    sigma=sigma, observed=y) All checks Accept
  3. 3.
    Diagnostic Checks & Guided Resampling We run NUTS on the accepted model and observe:
    • $\widehat{R}=1.2$ (too high),
    • 100 divergences.
    These failures reduce the Bayesian reliability score $B(M)$. RefineStat then resamples the likelihood or prior fragments (via the LLM) and retries inference until diagnostics ($\widehat{R}$, ESS, divergences, Pareto-$k$) fall within thresholds or the iteration limit is reached.
The final program $M^{*}$ converges ($\widehat{R}\approx 1$), shows zero divergences, and yields reliable ELPD-LOO. By pruning invalid distributions early and resampling based on diagnostic triggers, RefineStat iteratively refines candidates into a high-scoring $M^{*}$ with robust statistical reliability.
D.2 Examples of Structural Changes Introduced by RefineStat
In addition to refining a fixed model skeleton, RefineStat can introduce qualitatively new structural components during the resampling process. When the model families and diagnostic feedback indicates low reliability, RefineStat may alter the functional form of the model, add new parameters, or change the dependency structure to explore alternatives with higher Bayesian reliability scores.
Below, we present three programs generated for the same dataset and initialization. These examples illustrate how RefineStat’s refinement process can modify terms in the likelihood, expand the parameter space, or even change which variable is treated as observed.
Program 1: Standard linear regression alpha = pm.Normal("alpha", mu=0, sigma=10) beta = pm.Normal("beta", mu=0, sigma=10) sigma = pm.HalfNormal("sigma", sigma=1) mu = alpha + beta * year likelihood = pm.Normal("likelihood", mu=mu, sigma=sigma, observed=C)
Program 2: Nonlinear likelihood and an additional prior alpha = pm.Normal("alpha", mu=0, sigma=10) beta = pm.Normal("beta", mu=0, sigma=10) gamma = pm.Normal("gamma", mu=0, sigma=10) sigma = pm.HalfNormal("sigma", sigma=10) likelihood = (alpha + beta * year + gamma * year**2) * C + sigma * N observed = pm.Normal("observed", mu=likelihood, sigma=sigma, observed=C)
Program 3: Changing the response variable and dependency structure alpha = pm.Normal("alpha", mu=0, sigma=10) beta = pm.Normal("beta", mu=0, sigma=10) gamma = pm.Normal("gamma", mu=0, sigma=10) sigma = pm.HalfNormal("sigma", sigma=10) y_pred = alpha + beta * year + gamma * C y_obs = pm.Normal("y_obs", mu=y_pred, sigma=sigma, observed=N)
Summary of Structural Differences.
Program 1 uses a standard linear regression and models C as a linear function of year. Program 2 expands the model by introducing a new parameter (gammagamma) and applying a nonlinear transformation involving $\texttt{year}^{2}$, C, and N, leading to a different generative process. Program 3 changes the response variable entirely, modeling N instead of C, and modifies the dependency graph by incorporating C as a covariate in the linear predictor.
These examples show that RefineStat is capable of exploring alternative model families and introducing new structure, such as additional priors, nonlinear link functions, or altered observational targets whenever such modifications remain semantically valid.
Appendix E Prompt Design
E.1 Eight Schools
Description: A hierarchical model for the 8-schools data.
E.2 Dugongs
Description: A growth model for dugongs with missing data.
E.3 Surgical
E.4 GP
Description: Simulated data from a Poisson GP model.
E.5 Peregrine
Appendix F Experimental Setup
We set the convergence threshold to $\alpha_{R}=1.05$ for split-$\widehat{R}$, allow $\mathrm{ESS}_{\mathrm{bulk}}\geq\beta_{\mathrm{bulk}}=400$, and adopt a relaxed cutoff $\beta_{\mathrm{tail}}=100$ for $\mathrm{ESS}_{\mathrm{tail}}$ to accommodate lower sampling efficiency in the tails. For leave-one-out validation, $\widehat{\mathrm{elpd}}(M)$ must be finite, with at least $1-\epsilon=0.8$ of data points having Pareto shape values below $L_{\mathrm{cd}}=0.7$. These thresholds are used consistently when computing the reliability score across all models. We use Standard unconstrained generation as our baseline. Further, based on preliminary experiments we have chosen $\beta$ to be 4, $\alpha$ to be 2, and $R_{\max}$ as 100 for all experimental purposes.
We run experiments on a 48-core Intel Xeon Silver 4214R CPU with 2 NVidia RTX A5000 GPUs. RefineStat is implemented using PyTorch Paszke et al. (2019), and Itergen library Ugare et al. (2024a) for refining the parser-guided LLM generation infrastructure. We run all experiments for 10 seeds to reduce result randomness, and use a temperature range of 0.2 to 0.4.
Appendix G Error Analysis
G.1 Run rate failures
We categorized the failures by their root causes in different methods found during the run rate experiment:
  • Syncode eliminated many basic syntactic mistakes but still suffered semantic errors, such as incorrect distribution parameter names, type mismatches, referring to deprecated API functions (e.g., calling pm.sample_prior from an earlier PyMC release), and inventing non‐existent methods like pm.random_coefs.
  • RefineStat, in contrast, often produced models whose samplers failed to explore the correct posterior modes, leading to chains stuck in low‐density regions or divergent transitions, failures stemming from the model definitions rather than our decoding framework; RefineStat reduced both syntactic and semantic errors and avoided sampler pathologies by enforcing grammar and parameter validity during decoding.
G.2 Other Failure Modes
@@
Appendix H Ablation Study
H.1 Token Efficiency Analysis
To evaluate the computational cost associated with our framework, we measure the number of tokens consumed in generating a program under Itergen, Baseline (Unconstrained generation), RefineStat without Refinement Loop (RefineStat w/o RL), and RefineStat using Meta-LLama-3-8B. The token usage across five runs is recorded for each of these methods, from which we report the mean and standard deviation.
Dataset Itergen Baseline RefineStat w/o RL RefineStat Token
Mean Std Mean Std Mean Std Mean Std Ratio
Eight Schools $98.8$ $5.3$ $606.8$ $386.3$ $147.6$ $36.7$ $1503.0$ $409.5$ $2.5$
Dugongs $209.3$ $225.9$ $747.4$ $495.0$ $294.4$ $86.3$ $419.0$ $87.1$ $0.6$
GP $131.3$ $8.9$ $663.6$ $410.9$ $155.0$ $21.2$ $1660.0$ $253.5$ $2.5$
Peregrine $132.5$ $21.4$ $884.0$ $538.9$ $180.8$ $103.6$ $1740.0$ $418.0$ $2.0$
Surgical $97.3$ $4.3$ $624.4$ $260.6$ $113.2$ $9.2$ $1275.0$ $1348.1$ $2.0$
Average Token Ratio 1.9x
H.2 Memorization Effect
To further stress-test our approach against memorization, we performed two controlled prompt modifications across all datasets using Meta-Llama 3-8B.
Mean Std Mean Std Mean Std Mean Std Mean Std Mean Std